OpenStudy (anonymous):

how do i find the derivative of 4/(√x) ?

OpenStudy (anonymous):

y=4/x^1/2 y=4x^-1/2 dy/dx= 4* (-1/2)x^-3/2 dy/dx= -2x^-3/2

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

fan me pls thanx